Conchita noticed that there were fewer than 100 Hippos in the river. when she counted them by 2's. there was 1 hippo left over. in fact, when she counted them by 3's, 4's, 5's, or 6's there was always 1hippo left over.
how many Hippos were in the river??

Answer by venugopalramana(3286)

THE GIVEN CONDITION MEANS ,THE NUMBER OF HIPPOS LESS ONE IS DIVISIBLE BY 2,3,4,5,AND 6.SMALLEST UH NUMBER WOULD BE THE LCM OF THESE NUMBERS
2..|2,3,4,5,6
-----------------
..3|1,3,2,5,3
------------------
....1,1,2,5,1
HENCE LCM = 2*3*1*1*2*5*1=60
SO NUMBER OF GOATS = 60+1=61....NOTE THAT IT COULD BE
ANY MULTIPLE OF LCM=60 PLUS 1...THAT IS 121,181,ETC...